# MegaFusion: Extend Diffusion Models towards Higher-resolution Image Generation without Further Tuning

### v1

@@ 4.2 MegaFusion++
Dilated Convolution. Blurriness and semantic deviation in high-resolution images from diffusion models can be attributed to the constrained receptive field of UNet layers trained on fixed-resolution data, which lack comprehensive global information. For instance, as illustrated in Figure 2 (b), existing models trained on low-resolution images tend to synthesize multiple rabbits in different local areas due to insufficient receptive fields, leading to semantic inaccuracies. Inspired by ScaleCrafter [11], we modify the convolutional kernels of the UNet-based denoiser $\epsilon_{\theta}$ to incorporate dilated convolutions [49] with a specific dilation rate $\delta$. This broadens the model’s receptive field without additional tuning, allowing for better incorporation of global information.
@@
For the sake of simplicity, we omit channel dimension and convolution biases here, focusing on the modification of weight parameters to transform standard convolutions into dilated counterparts. Given a feature map $F\in\mathbb{R}^{m\times n}$ and a convolutional kernel $k\in\mathbb{R}^{r\times r}$, the standard convolution can be represented as: $(F*k)(p)=\sum_{s+t=p}F(s)\cdot k(t)$. In contrast, the corresponding dilated convolution, with dilation rate $\delta$, can be expressed as: $(F*_{\delta}k)(p)=\sum_{s+\delta t=p}F(s)\cdot k(t)$. Here, $p,k,$ and $t$ denote spatial locations within the feature map and convolution kernel, respectively.
@@
Rather than replacing all convolutions with dilated ones, which may lead to catastrophic quality decline, we selectively apply this modification to the middle layers of UNet. The insight here is that: we wish to broaden the receptive field in the bottleneck to aggregate global information, while simultaneously preserving original priors at higher resolution to sample nearby features for enhancing details.
@@
Noise Re-scheduling. Similar to discoveries in simple diffusion [18] and relay diffusion [41], we observe that identical noise levels impact images differently across various resolutions, as depicted in Figure 2 (c), resulting in varying signal-to-noise ratios (SNR) at the same timestep.
@@
According to the SNR definition in previous work [18]: $SNR_{t}=\frac{(\sqrt{\bar{\alpha}}_{t})^{2}}{(\sqrt{1-\bar{\alpha}}_{t})^{2}}=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}$. Considering a low-resolution image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$ and a high-resolution one $\mathbf{x}^{\prime}\in\mathbb{R}^{3\times H^{\prime}\times W^{\prime}}$ with $H^{\prime}=2H$ and $W^{\prime}=2W$, if we downsample $\mathbf{x}^{\prime}$ to ${\mathbf{x}^{\prime}}_{down}\in\mathbb{R}^{3\times H\times W}$, the SNR at timestep $t$ of ${\mathbf{x}^{\prime}}_{down}$ (denoted as ${SNR}_{down}^{H^{\prime}\times W^{\prime}}$) in comparison to $\mathbf{x}$ (represented as ${SNR}^{H\times W}$) will exhibit the following relationship: ${SNR}^{H\times W}=\gamma\cdot{SNR}_{down}^{H^{\prime}\times W^{\prime}}$.
@@
Assuming the original noise scheduler at $H\times W$ is denoted as $\bar{\alpha}_{t}$, the revised scheduler ${\bar{\alpha}_{t}}^{\prime}$ at higher resolution $H^{\prime}\times W^{\prime}$ should satisfy: $\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}=\gamma\cdot\frac{{\bar{\alpha}_{t}}^{\prime}}{1-{\bar{\alpha}_{t}}^{\prime}}$. This yields the relationship: ${\bar{\alpha}_{t}}^{\prime}=\frac{{\bar{\alpha}_{t}}}{\gamma-(\gamma-1){\bar{\alpha}_{t}}}$. Incorporating this into the high-resolution noise scheduler initialization gives a new ${\bar{\alpha}_{t}}’$ ’ sequence. This process, called noise re-scheduling, adjusts noise levels to better suit higher-resolution image generation, thereby improving synthesis quality and fidelity.
@@ 4.3 Further Application on other Models
Pixel-space Diffusion Models. Similar to latent-space models, our proposed MegaFusion is equally applicable to pixel-space diffusion models, e.g. Floyd [6]. The primary distinction lies in performing the truncate and relay operation directly in pixel space. This implies that Equations 2, 3 and 4 can be adapted as follows:
@@
Diffusion Models with Extra Conditions. Beyond the scope of text-to-image foundation models, our methodology can extend to diffusion models that incorporate extra input conditions, such as ControlNet [51] and IP-Adapter [48]. These models use both text condition $c_{T}$ and image condition $c_{I}$ as input. Consequently, Equation 1 can be reformulated to accommodate both conditions:
$\mathbf{z}_{t}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon(\mathbf{z_{t}},t,c_{T},c_{I}))+\sigma_{t}\epsilon$ (7)
@@ 5.1 Experiment Settings
Implementation Details. We evaluate text-to-image diffusion models in both latent space (SDM 1.5 [35] and SDXL [30]) and pixel space (Floyd [6]). All models use DDIM [40] for $T=50$ steps of sampling except in cases explicitly stated otherwise. Given that SDM is trained with a fixed resolution of $512\times 512$, we choose to generate high-resolution images of $1024\times 1024$ for quantitative comparison. Specifically, we orchestrate generation across $k=3$ resolutions: $512$, $768$ and $1024$, with respective denoising steps of $T_{1}=40$, $T_{2}=5$, and $T_{3}=5$. Furthermore, our proposed MegaFusion can be applied to synthesize $2048\times 2048$ and even higher-resolution images with SDM for qualitative assessment. SDXL defaults to synthesizing images of $1024\times 1024$, considering the balance of computational costs, we discard the Refiner module and employ two-stage generation, $1024$ and $2048$, with their respective denoising steps being $T_{1}=40$ and $T_{2}=10$.
On the other hand, Floyd, a 3-stage cascaded model, sequentially upscales images from $64\times 64$ to $256\times 256$, culminating in $1024\times 1024$ images. Due to computational constraints, only the initial two stages of Floyd are employed in our experiments. The first stage necessitates $100$ steps of DDIM sampling ($T_{1}=80$ for generating $64\times 64$ images, and $T_{2}=20$ for $128\times 128$), while the second stage requires $50$ steps ($40$ for $T_{1}=256\times 256$ and $T_{2}=10$ for $512\times 512$).
Bicubic upsampling serves as the default non-parametric Upsampler $\Phi$. For dilated convolutions, we set the dilation rate $\delta=2$, and select the hyperparameter $\gamma=4$ for noise re-scheduling. All experiments are conducted on a single Nvidia RTX A40 GPU, with SDM and SDXL in $float16$ half-precision, and Floyd at in $float32$ full-precision.
Evaluation Datasets. We assess our method and baseline models on the MS-COCO [24] dataset, which comprises approximately 120K images in total, each accompanied by 5 captions. Due to the computational costs of high-resolution generation, we randomly sample 10K images from MS-COCO, assigning a fixed caption to each as input. To ensure a consistent comparison, we utilize the same random seed for each image across different methods, neutralizing randomness. For qualitative human evaluations, we use commonly available prompts from the Internet as text conditions, and conditional images provided by the original code repositories as extra inputs for IP-Adapter and ControlNet.
@@
Figure 4: Qualitative results of incorporating our method to models with extra conditional inputs. Our proposed MegaFusion can be universally applied across various diffusion models, providing the capability for higher-resolution image generation with better semantics and fidelity. Please refer to the Appendix for more qualitative results.
@@ 5.2 Quantitative Results
Objective Metrics. We compare the performance of both latent-space and pixel-space diffusion models boosted by our MegaFusion with their baseline counterparts on the MS-COCO [24] dataset. Here, [model-MegaFusion] refers to models employing truncate and relay strategy to bridge multi-resolution generation, while [model-MegaFusion++] denotes advanced models that incorporate dilated convolution and noise re-scheduling. We also compare with several existing state-of-the-art methods, such as SDM with StableSR [44], ScaleCrafter [11], and DemoFusion [7], which are limited to specific latent-space models and less efficient.
The results in Table 1 highlight significant improvements with MegaFusion across all metrics, including image quality, semantic accuracy, and especially computational efficiency. This indicates that our MegaFusion effectively extends the generative capabilities of existing diffusion models towards synthesizing high-resolution images with correct semantics and details while using only about 40% of the computational cost. Furthermore, incorporating dilated convolution and noise re-scheduling further improves performance on metrics such as $\mathrm{FID}_{r}$, $\mathrm{KID}_{r}$, CLIP-T, and ROUGE, reflecting an improvement in generation diversity and alignment with real images and text conditions.
Human Assessment. To complement our objective analysis, we conduct a human-centric evaluation focusing on image quality and semantic integrity. Concretely, utilizing the same text and random seed as inputs, we synthesize higher-resolution images via standard models (SDM and Floyd) and their MegaFusion-boosted counterparts. Participants are asked to rate the outputs with a score from 1 to 5 (the higher, the better), considering both image quality and semantic accuracy. Moreover, they also need to select their favourite image among the options for preference rating.
The results in Table 2 affirm that our MegaFusion significantly improves the performance of higher-resolution image generation in terms of image quality and semantic accuracy. Additionally, our advanced MegaFusion++ demonstrates the potential for greater improvements. This evidence underscores MegaFusion’s ability to elevate pre-trained models, enabling them to produce higher-resolution images with superior quality and accurate semantics.
@@ 5.3 Qualitative Results
Comparison on models with additional conditions. We further apply MegaFusion to diffusion models equipped with extra input conditions, such as IP-Adapter and ControlNet, as illustrated in Figure 4. Our MegaFusion exhibits universal applicability, significantly extending the capacity of various diffusion models to synthesize high-quality images of higher resolutions, which not only adhere to the input conditions but also maintain semantic integrity.
@@
Table 3: Ablation Study on proposed modules in MegaFusion on MS-COCO [24]. The modules gradually improve the higher-resolution image generation quality of existing diffusion models, especially in comparison with real images.
@@ 5.4 Ablation Studies
Proposed strategy & modules. To evaluate the efficacy of our proposed strategy and components, we assess several model variants in both latent and pixel spaces. Here, ‘T&R’, ‘D’, and ‘R’ represent truncate and relay strategy, dilated convolution, and noise re-scheduling, respectively. The results in Table 3 demonstrate that our strategy and modules substantially elevate the quality and diversity of contents generated by foundational generative models such as SDM ($1024\times 1024$) and Floyd ($128\times 128$), especially improving the quality and fidelity to real-world images.
Upsampler $\Phi$. In our truncate and relay strategy, a non-parametric Upsampler plays a crucial role in bridging generation processes across different resolutions. To discern the optimal choice, we evaluate several configurations of the SDM-MegaFusion++ model on MS-COCO dataset. These variants include Config-A with bilinear upsampling; Config-B with bicubic upsampling; Config-C incorporating bicubic coupled with a $5\times 5$ Gaussian filter, and Config-D employing bicubic alongside a $3\times 3$ edge-enhancement kernel. As depicted in Table 4, SDM-MegaFusion++ with Config-B emerges as the superior method in terms of both FID and KID metrics. Consequently, we have designated bicubic upsampling as the default choice.
@@ 6 Conclusion
In this paper, we introduce MegaFusion, a tuning-free approach designed to tackle the challenges of synthesizing higher-resolution images, which effectively addresses issues of semantic inaccuracies and object replication. Our method adopts an innovative truncate and relay strategy to elegantly connect generation processes across different resolutions, efficiently synthesizing high-quality higher-resolution images of megapixels and various aspect ratios in a coarse-to-fine manner. By incorporating orthogonal dilated convolutions and noise re-scheduling, we further adapt the model priors towards higher resolution. The versatility and efficacy of our MegaFusion enable it universally applicable to both latent-space and pixel-space diffusion models, as well as their extensions with extra conditions. Extensive experiments have validated the superiority of MegaFusion, demonstrating its capability to generate images at higher resolutions with only about 40% of the original computational cost.
@@ Appendix A Implementation Details
More Details on Floyd-MegaFusion. We have evaluated the higher-resolution image generation performance of the Floyd [6] model at resolutions of $128\times 128$ and $512\times 512$. For $128\times 128$ resolution, we directly apply MegaFusion to the first stage of the Floyd model. As for the comparison at $512\times 512$ resolution, we utilize the first two stages of the Floyd model. Considering that the quality of the results from the first stage generation would significantly affect the second generation stage, we opt for using the $64\times 64$ images generated by the original first stage model as inputs of both the baseline and our boosted Floyd-MegaFusion. That is, higher-resolution image generation is only performed under the second generation stage. Ultimately, the experimental results presented in Table 1 effectively demonstrate the universality and effectiveness of our proposed MegaFusion. Furthermore, we also conduct experiments where $128\times 128$ out-of-distribution images are generated in the first stage, followed by $512\times 512$ resolution images in the second stage. This further demonstrates that MegaFusion maintains semantic accuracy across all stages of generation.
Details on Human Evaluation. To more effectively reflect the performance of different models in generating high-resolution images, we have recruited 10 volunteers with a background in image generation research for human evaluation. Specifically, the evaluators are asked to follow these rules: (i) Rate unknown source images on a score from 1 to 5 for both image quality and semantic accuracy, with higher scores indicating better quality; and (ii) Observe the results generated by different models with the same input conditions and select their favourite one based on a comprehensive consideration of quality and semantics.

### v2

@@ 4.2 MegaFusion++
Dilated Convolution. Blurriness and semantic deviation in high-resolution images generated by diffusion models often stem from the constrained receptive field of UNet layers trained on fixed-resolution data, lacking comprehensive global context. As illustrated in Figure 2 (b), existing models trained on low-resolution images tend to synthesize multiple rabbits in different local regions due to insufficient receptive fields, leading to semantic inaccuracies. Inspired by ScaleCrafter [14], we modify the convolutional kernels of the UNet-based denoiser $\epsilon_{\theta}$ to incorporate dilated convolutions [52] with a specific dilation rate $\delta$. This broadens the receptive field without additional tuning, allowing for better incorporation of global information.
@@
For the sake of simplicity, we omit channel dimensions and convolution biases here, focusing on modifying weight parameters to transform standard convolutions into dilated ones. Given a feature map $F\in\mathbb{R}^{m\times n}$ and a convolutional kernel $k\in\mathbb{R}^{r\times r}$, the standard convolution can be represented as: $(F*k)(p)=\sum_{s+t=p}F(s)\cdot k(t)$. In contrast, the corresponding dilated convolution, with dilation rate $\delta$, can be expressed as: $(F*_{\delta}k)(p)=\sum_{s+\delta t=p}F(s)\cdot k(t)$. Here, $p,k,$ and $t$ denote spatial locations within the feature map and convolution kernel, respectively.
@@
Following previous practices [14], instead of replacing all convolutions with dilated ones, which may lead to catastrophic quality decline, we selectively apply this modification to the middle layers of UNet. The insight here is that: we broaden receptive fields in the bottleneck to aggregate global context, while preserving original priors at higher resolution to sample nearby features for enhancing details.
@@
Noise Re-scheduling. Consistent with discoveries in simple diffusion [21] and relay diffusion [43], we observe that identical noise levels impact images differently across various resolutions, as illustrated in Figure 2 (c), leading to varying signal-to-noise ratios (SNR) at the same timestep.
@@
According to the SNR definition in previous work [21]: $SNR_{t}=\frac{(\sqrt{\bar{\alpha}}_{t})^{2}}{(\sqrt{1-\bar{\alpha}}_{t})^{2}}=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}$. Given a low-resolution image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$ and a high-resolution one $\mathbf{x}^{\prime}\in\mathbb{R}^{3\times H^{\prime}\times W^{\prime}}$ with $H^{\prime}>H$ and $W^{\prime}>W$, if we downsample $\mathbf{x}^{\prime}$ to ${\mathbf{x}^{\prime}}_{down}\in\mathbb{R}^{3\times H\times W}$, the SNR at timestep $t$ of ${\mathbf{x}^{\prime}}_{down}$ (denoted as ${SNR}_{down}^{H^{\prime}\times W^{\prime}}$) in comparison to $\mathbf{x}$ (represented as ${SNR}^{H\times W}$) will exhibit the following relationship: ${SNR}^{H\times W}=\gamma\cdot{SNR}_{down}^{H^{\prime}\times W^{\prime}}$.
@@
Assuming the original noise scheduler at $H\times W$ is denoted as $\bar{\alpha}_{t}$, the revised scheduler ${\bar{\alpha}_{t}}^{\prime}$ at higher resolution $H^{\prime}\times W^{\prime}$ should satisfy: $\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}=\gamma\cdot\frac{{\bar{\alpha}_{t}}^{\prime}}{1-{\bar{\alpha}_{t}}^{\prime}}$. This yields the relationship: ${\bar{\alpha}_{t}}^{\prime}=\frac{{\bar{\alpha}_{t}}}{\gamma-(\gamma-1){\bar{\alpha}_{t}}}$. Incorporating this into the high-resolution noise scheduler initialization gives a new ${\bar{\alpha}_{t}}$ sequence. This process, termed noise re-scheduling, adjusts noise levels to better suit higher-resolution image generation, thereby improving synthesis quality and fidelity.
@@ 4.3 Further Application on other Models
Pixel-space Diffusion Models. MegaFusion is equally applicable to pixel-space diffusion models, such as Floyd [8], with the primary difference being that the truncate and relay operation is performed directly in pixel space. Consequently, Equations 2, 3, and 4 are adapted as follows:
@@
Diffusion Models with Extra Conditions. Our methodology can also extend to diffusion models that incorporate extra input conditions, such as ControlNet [54] and IP-Adapter [51]. These models utilize both text condition $c_{T}$ and image condition $c_{I}$ as inputs. Consequently, Equation 1 can be reformulated to accommodate both conditions:
$\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon(\mathbf{z}_{t},t,c_{T},c_{I}))+\sigma_{t}\epsilon$ (7)
@@ 5.1 Experiment Settings
Implementation Details. We evaluate text-to-image diffusion models in both latent space (SDM 1.5 [37], SDXL [32] and SD3 [10]) and pixel space (Floyd [8]). All models use DDIM [42] for $T=50$ steps of sampling unless explicitly stated otherwise. Given that SDM is trained with a fixed resolution of $512\times 512$, we choose to generate high-resolution images of $1024\times 1024$ for quantitative comparison. Specifically, we orchestrate generation across $k=3$ resolutions: $512$, $768$ and $1024$, with respective denoising steps of $T_{1}=40$, $T_{2}=5$, and $T_{3}=5$. Furthermore, our proposed MegaFusion can be applied to synthesize even higher-resolution images with SDM for qualitative assessment. SDXL defaults to synthesizing images of $1024\times 1024$, considering the balance of computational costs, we discard the Refiner module and employ two-stage generation, $1024$ and $2048$, with their denoising steps being $T_{1}=40$ and $T_{2}=10$. For SD3, which defaults to denoise 28 steps for generating $1024\times 1024$ images, we iterate $20$ steps at $1024$ resolution and $8$ steps at $2048$ resolution.
On the other hand, Floyd, a 3-stage cascaded model, sequentially upscales images from $64\times 64$ to $256\times 256$, culminating in $1024\times 1024$ images. Due to computational constraints, only the first two stages of Floyd are employed in our experiments. The first stage necessitates $100$ sampling steps ($T_{1}=80$ for generating $64\times 64$ images, and $T_{2}=20$ for $128\times 128$), while the second stage requires $50$ steps ($T_{1}=40$ for $256\times 256$ and $T_{2}=10$ for $512\times 512$).
Bicubic upsampling serves as the default non-parametric Upsampler $\Phi$. For typical 2×2$\times$2 × higher-resolution generation, we set the dilation rate $\delta=2$, and select the hyperparameter $\gamma=4$ for noise re-scheduling. For classifier-free guidance, to ensure a fair comparison, we apply the default weight $w$ of official implementations across all methods: $w=7.0$ for SDM, SDXL, SD3, and Floyd-Stage 1, and $w=4.0$ for Floyd-Stage 2. All experiments are conducted on a single Nvidia RTX A40 GPU, with SDM, SDXL, and SD3 in $float16$ precision, and Floyd at in $float32$ precision.
Evaluation Datasets. We assess our method and baseline models on the MS-COCO [26] dataset, which comprises approximately 120K images in total, each accompanied by 5 captions. Due to the computational costs of high-resolution generation, we randomly sample 10K images from MS-COCO, assigning a fixed caption to each as input. To ensure consistent comparisons, we utilize the same random seed for each image across methods, neutralizing randomness. For qualitative human evaluations, we use commonly available prompts from the Internet as text conditions and conditional images from the official code repositories as extra inputs for IP-Adapter and ControlNet.
@@
Figure 4: Qualitative results of incorporating MegaFusion to models with extra conditional inputs. MegaFusion can be universally applied across various diffusion models, providing the capability for higher-resolution image generation with better semantics and fidelity.
@@ 5.2 Quantitative Results
Objective Metrics. We evaluate the performance of both latent-space and pixel-space diffusion models boosted by MegaFusion against their baseline counterparts on the MS-COCO [26] dataset. Here, [model-MegaFusion] refers to models employing truncate and relay strategy to bridge multi-resolution generation, while [model-MegaFusion++] denotes advanced models incorporating dilated convolution and noise re-scheduling. We also compare with several existing state-of-the-art methods, such as SDM and SDXL with StableSR [46] or RealESRGAN [47], ResAdapter [7], ScaleCrafter [14], and DemoFusion [9], which are limited to specific latent-space models and less efficient.
The results in Table 1 highlight significant improvements with MegaFusion across all metrics, including image quality, semantic accuracy, and especially computational efficiency. This confirms that MegaFusion effectively extends the generative capabilities of existing diffusion models towards synthesizing high-resolution images with correct semantics and details at only about 40% of the original computational cost. Moreover, incorporating dilated convolution and noise re-scheduling further improves performance on metrics such as $\mathrm{FID}_{r}$, $\mathrm{KID}_{r}$, CLIP-T, and ROUGE, reflecting improved generation diversity and better alignment with real images and text conditions.
Human Assessment. To complement our objective analysis, we conduct a human-centric evaluation focusing on image quality and semantic integrity. Concretely, utilizing identical text prompts and random seed, we synthesize higher-resolution images via standard models (SDM and Floyd) and their MegaFusion-boosted counterparts. Participants are asked to rate the outputs with a score from 1 to 5 (higher is better), considering both image quality and semantic accuracy. Additionally, they also need to select their preferred image among the options for preference rating.
The results in Table 2 affirm that our MegaFusion significantly improves the performance of higher-resolution image generation in terms of image quality and semantic accuracy. Additionally, our advanced MegaFusion++ further demonstrates potential for even greater improvements. This evidence underscores MegaFusion’s ability to elevate pre-trained models, enabling them to produce higher-resolution images with superior quality and precise semantics.
@@ 5.3 Qualitative Results
Comparison on models with additional conditions. We further apply MegaFusion to diffusion models equipped with extra input conditions, such as IP-Adapter and ControlNet, as illustrated in Figure 4. Our MegaFusion exhibits universal applicability, significantly extending the capacity of various diffusion models to synthesize high-quality images of higher resolutions, which not only adhere to the input conditions but also maintain semantic integrity. Please refer to the Appendix for more qualitative results.
@@
Table 3: Ablation Study on proposed modules in MegaFusion on MS-COCO. The modules gradually improve the higher-resolution generation quality, especially in comparison with real images.
@@ 5.4 Ablation Studies
Proposed strategy & modules. To evaluate the efficacy of our proposed strategy and components, we assess several model variants in both latent and pixel spaces. Here, ‘T&R’, ‘D’, and ‘R’ represent the truncate and relay strategy, dilated convolution, and noise re-scheduling, respectively. The results in Table 3 demonstrate that our strategy and modules significantly elevate the quality and diversity of contents generated by generative models such as SDM ($1024\times 1024$) and Floyd ($128\times 128$), especially in improving the quality and fidelity to real-world images.
Upsampler $\Phi$. In our truncate and relay strategy, a non-parametric Upsampler is crucial in bridging generation processes across different resolutions. To determine the optimal choice, we evaluate several variants of the SDM-MegaFusion++ model on MS-COCO dataset. These include Config-A with bilinear upsampling; Config-B with bicubic upsampling; Config-C incorporating bicubic coupled with a $5\times 5$ Gaussian filter, and Config-D employing bicubic alongside a $3\times 3$ edge-enhancement kernel. As depicted in Table 4, SDM-MegaFusion++ with Config-B outperforms others in terms of both FID and KID metrics, leading us to adopt bicubic upsampling as the default choice.
@@ 6 Conclusion
In this paper, we introduce MegaFusion, a tuning-free approach designed to tackle the challenges of synthesizing higher-resolution images, effectively resolving issues of semantic inaccuracies and object replication. Our method adopts an innovative truncate and relay strategy to elegantly connect generation processes across different resolutions, synthesizing higher-resolution images with megapixels and various aspect ratios in a coarse-to-fine manner. By orthogonally incorporating dilated convolutions and noise re-scheduling, we further adapt model priors towards higher resolution. The versatility and efficacy of MegaFusion enable it universally applicable to both latent-space and pixel-space diffusion models, as well as their extensions with extra conditions. Extensive experiments have validated the superiority of MegaFusion, demonstrating its capability to generate higher-resolution images with approximately 40% of the original computational cost.
@@ Appendix A Implementation Details
More Details on Floyd-MegaFusion. We have evaluated the higher-resolution image generation performance of Floyd [8] at resolutions of $128\times 128$ and $512\times 512$. For $128\times 128$ resolution, we directly apply MegaFusion to the first stage of Floyd. As for the comparison at $512\times 512$ resolution, we utilize the first two stages of Floyd. Considering that the quality of the results from the first stage generation would significantly affect the second generation stage, we opt for using the $64\times 64$ images generated by the original first stage model as inputs of both the baseline and our boosted Floyd-MegaFusion. That is, higher-resolution image generation is only performed under the second generation stage. Ultimately, the experimental results presented in Table 1 of our submitted manuscript effectively demonstrate the universality and effectiveness of our proposed MegaFusion. Furthermore, we also conduct experiments where $128\times 128$ out-of-distribution images are generated in the first stage, followed by $512\times 512$ resolution images in the second stage. This further demonstrates that MegaFusion maintains semantic accuracy across all stages of generation.
Details on Human Evaluation. To more effectively reflect the performance of different models in generating high-resolution images, we have recruited 10 volunteers with a background in image generation research for human evaluation. Specifically, the evaluators are asked to follow these rules: (i) Rate unknown source images on a score from 1 to 5 for both image quality and semantic accuracy, with higher scores indicating better quality; and (ii) Observe the results generated by different models with the same input conditions and select their favourite one based on overall quality and semantic accuracy.